Materials used by the Instrument Division of XPort Industries are currently purchased from outside suppliers at a cost of $229 per unit. However, the same materials are available from the Components Division. The Components Division has unused capacity and can produce the materials needed by the Instrument Division at a variable cost of $190 per unit.

Assume that a transfer price of $218 has been established and that 45,600 units of materials are transferred, with no reduction in the Components Division’s current sales.

a. How much would XPort Industries’ total income from operations increase?

b. How much would the Instrument Division’s income from operations increase?

c. How much would the Components Division’s income from operations increase?
Answer #1

a. XPort Industries’ total income from operations increase by = ($229 - $190)*45600 = $1778400

b. Instrument Division’s income from operations increase by = (229-218)*45600 = $501,600

c. Components Division's income from operations increase by = (218-190)*45600 = $1276800